OpenStudy (anonymous):

elimination 6x - 3y = 15 7x + 4y = 10

OpenStudy (anonymous):

Start by multiplying top equation by 4, and second equation by 3. Then you can eliminate y :D

OpenStudy (anonymous):

Thats how I would start :)

OpenStudy (anonymous):

so what i do next

OpenStudy (anonymous):

Did you get rid of y? So what do you have left?

OpenStudy (anonymous):

i left with x and = but i dont know if to add or subtact

OpenStudy (anonymous):

|24x - 12y = 60 21x + 12y = 30 and i cancel the y already so dont worry bout the y so what i do with x and =

OpenStudy (anonymous):

Well, add up the x's and numbers to get: 45x = 90 So x = 2 You have x now :D YAY Now plug this into one of your two original equations: Lets say 6x - 3y = 15 and solve for y So you have: -3y = -6x+15 3y=6x-15 y=2x-5 So plug x = 2 into (y=2x-5)

OpenStudy (anonymous):

Wow....Zelda.... you could be a teacher :)

OpenStudy (anonymous):

@kelliegirl33, thanks, but I'm a Law student (3rd year in Cornell) So not a teacher :P

OpenStudy (anonymous):

how i solve for y

OpenStudy (anonymous):

I told you. We found x, put it in for one of the two equations and solve for Y.

OpenStudy (anonymous):

my answer is (2,1)

OpenStudy (anonymous):

Close, y = -1 :p Anyhow The solution should be (2,-1)

OpenStudy (anonymous):

thank u vey much @Zelda

OpenStudy (anonymous):

@lovechama, no prob kiddo.